A train running at the speed of 60 km/hr crosses a pole in 9 seconds. What is the length of the train?
solution:
Speed= 60X5/18 m/sec = 50/3m/sec
length of the train = S X D = 50/3 X 9m = 150m
when do we use 5/18 ?

Answer by mananth(16949)

1 km/hour = 1000 meters/3600 seconds
=5/18 meters / second
..
60 km/hour = 60 *5/18 = 16.66 m/second